OpenStudy (anonymous):

MEDAL REWARDED Perform the requested operation or operations. f(x) = x-8/7 ; g(x) = 7x + 8, find g(f(x)).

OpenStudy (anonymous):

@John_ES

OpenStudy (anonymous):

Answer choices: g(f(x)) = x + 16 g(f(x)) = 7x + 48 g(f(x)) = x - 8/7 g(f(x)) = x

OpenStudy (john_es):

As you see, you only need to plug in the f(x) into the g(x).

OpenStudy (john_es):

\[g(f(x))=7f(x)+8=7(x−8)/7+8=x\]

OpenStudy (anonymous):

@John_ES so it just equals x correct?

OpenStudy (john_es):

True.